Six different national brands of chocolate chip cookies were randomly selected at the supermarket. The grams of fat per serving are as follows: 9; 9; 11; 6; 8; 8. Assume the underlying distribution is approximately normal.

(iii) Calculate the error bound. (Round your answer to two decimal places.)
Answer #1

sample mean, xbar = 8.5
sample standard deviation, s = 1.6432
sample size, n = 6
degrees of freedom, df = n - 1 = 5

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 2.571


ME = tc * s/sqrt(n)
ME = 2.571 * 1.6432/sqrt(6)
ME = 1.72

Error bound = 1.72

### Here, confidence interval is not given , so i consider 95% . If is there any changes plz let me know